# Coherent Local Explanations for Mathematical Optimization

### v1

@@ 3 Methodology
  1. 1.
    The function $\bm{x}\mapsto f(\bm{x};\bm{\theta})$ is affine.
  2. 2.
    The function $\bm{x}\mapsto\delta(\bm{x},\mathbb{X}(\bm{\theta}))$ is convex.
The proof of this proposition follows from applying composition rules of convex functions (Boyd & Vandenberghe, 2004). When the conditions in this proposition are satisfied, every local minimum of the optimization problem (10) is a global minimum. We can use first-order methods to find such a minimum if the functions are differentiable. We note that for $\delta$ as defined in (9) we have that (8) is convex in $\bm{\beta}$ if the functions $\bm{x}\mapsto\gamma_{t}(\bm{x},\bm{\theta}^{i})$ are convex in $\bm{x}$ for $t\in[T]_{1}$ since $x\mapsto\max\{0,x\}$ is convex and nondecreasing in $x$.
Weights.
We define the weights (similarly as in LIME) as the radial basis function kernel
$w^{i}=\exp(-d(\bm{\theta}^{i},\bm{\theta}^{0})^{2}/\nu^{2}),\quad i\in[N]_{0},$ (11)
where $\nu$ is the kernel parameter and $d$ is a distance function.
Sampling.
We recall that contrary to ML models, optimization models do not require model training per se. Therefore, $\bm{\theta}^{i}$ cannot be sampled according to the train data distribution. Depending on the context, $\bm{\theta}^{i}$ can be sampled from relevant distributions, or with pre-determined rules, e.g., discretization.
Besides, we note that for some values of $\bm{\theta}^{i}$ the optimization problem might be infeasible or unbounded. We therefore ensure the generated dataset $\mathcal{D}$ contains only feasible, bounded instances of the optimization problem.
Binary Components.
We opt for logistic regressors as interpretable surrogate models for the output components of the optimization problem that are restricted to binary values. Let $\mathcal{B}\subseteq\{f,x_{1},\ldots,x_{p}\}$ be the set of the binary components. Then for a binary component $c\in\mathcal{B}$ we consider predictors of the form $g_{c}(\bm{\theta})=\sigma(\bm{\beta^{\intercal}}_{c}\bm{\theta})$ with $\sigma:\mathbb{R}\to[0,1]$ the sigmoid function. The relative values of vector $\bm{\beta}_{c}$ then tell the user the feature importance for the probability of the component being 00 or $1$. For the binary components, we measure the prediction accuracy using the log-loss. The total accuracy loss then becomes
$\ell_{A}(\bm{\beta},h(\bm{\theta}^{i}))=$ $\lambda_{A_{1}}\sum_{c\in\overline{\mathcal{B}}}\|\bm{\beta^{\intercal}}_{c}\bm{\theta}^{i}-h_{c}(\bm{\theta}^{i})\|^{2}$ (12)
$-\lambda_{A_{2}}\sum_{c\in\mathcal{B}}h_{c}(\bm{\theta}^{i})\ln(\sigma(\bm{\beta^{\intercal}}_{c}\bm{\theta}^{i}))$
$\qquad\qquad+(1-h_{c}(\bm{\theta}^{i}))\ln(1-\sigma(\bm{\beta^{\intercal}}_{c}\bm{\theta}^{i})),$
@@
Guaranteed Objective Coherence in the Linear Case.
Assume our present problem (2) has a linear objective function, i.e., it is of the form
$\min$ $\bm{\hat{c}}^{\intercal}\bm{x}$
$s.t.\quad$ $\bm{x}\in\mathbb{X}(\bm{\theta})$
and assume that only the feasible region is sensitive, i.e., $\bm{\hat{c}}$ remains fixed. In this case, we can fit $p+1$ independent linear models for each component in $c\in\{f,x_{1},\ldots,x_{p}\}$ by solving the classical weighted mean-square problem
$\min_{\bm{\beta}_{c}}\ \sum_{i=0}^{N}w^{i}\|\bm{\beta}_{c}^{\intercal}\bm{\theta}^{i}-h_{c}(\bm{\theta}^{i})\|^{2}.$
If the minimizers are unique, the corresponding linear predictors provably fulfill the coherence condition (3), i.e., in this case we do not need to apply the regularizer $R_{C}$ to achieve coherence condition (3). However, it may happen that condition (4) is violated as the example from the introduction shows. A proof of the latter coherence statement can be found in appendix A.3 (Theorem A.1).
@@ 4 Experiments
In this section, we present three experiments. Each experiment considers a distinct optimization model and solver. The first experiment is used as a proof of concept, where we will show that CLEMO approximates the optimal decision and objective value function well for an instance of the Shortest Path problem (SPP) with a single sensitive parameter. For a second experiment, in an extensive study we consider exact solutions of various instances of the Knapsack problem (KP). We compare the quality of explanations found by CLEMO to benchmarks by analyzing local accuracy, coherence, and stability of the found explanations when subjected to resampling. Lastly, we generate explanations for the Google OR-Tools heuristic (Furnon & Perron, ) applied to an instance of the Capacitated Vehicle Routing problem (CVRP). The code of our experiments can be found at https://github.com/daanotto/CLEMO.
4.1 Setup
Unless stated otherwise, all upcoming experiments use the following setup. Given an optimization problem for a given parameter vector $\bm{\theta}^{0}$, we create a training data set $\mathcal{D}$ of size 1000 by sampling $\bm{\theta}^{i}\sim\mathcal{N}(\bm{\theta}^{0},0.2\bm{\theta}^{0})$. The sample’s proximity weights $w^{i}$ are determined using the rbf kernel (11) with Euclidean distance and parameter $\nu$ equal to the mean distance to $\bm{\theta}^{0}$ over the training data set $\mathcal{D}$.
As a benchmark for CLEMO we consider generating explanations with the LIME-type method described in Section 3. We solve problem (5) where we fit logistic regressors for all binary output components and linear models for all other output components without any complexity regularization. We refer to this benchmark as LR. For CLEMO we use the loss function stated in (10), where $\ell_{A}$ is given as in (12) and $R_{C}$ is given as in (8) with $\delta$ defined as in (9). This way we can compare CLEMO to the benchmark on local accuracy (12) and incoherence (8). In CLEMO, each term of the total loss function is weighted with hyperparameters $\lambda_{A_{1}},\lambda_{A_{2}},\lambda_{C_{1}},$ and $\lambda_{C_{2}}$ as
$\lambda_{A_{1}}\ell_{A_{1}}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))$ $+\lambda_{A_{2}}\ell_{A_{2}}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))$
$+\lambda_{C_{1}}R_{C_{1}}(g(\bm{\theta}^{i}))$ $+\lambda_{C_{2}}R_{C_{2}}(g(\bm{\theta}^{i})).$
@@ 4.2 Shortest Path Problem
As a first experiment, we explain an instance of the Shortest Path where possible cost-changes depend on one single parameter. An instance of the SPP is given by a connected graph $G=(V,E,\bm{c})$, with nodes $V$, edges $E$ and edge-costs $\bm{c}$, and specified start and terminal nodes $s,t\in V$. The objective is to find a path between $s$ and $t$ of minimum costs. Several methods exist for solving the SPP in polynomial time (Gallo & Pallottino, 1988). Here, we use Dijkstra’s algorithm.
@@
value
Decision
vector
Objective
Feasible
region
@@
We study the parametric version of the shortest path problem (SPP-$\theta$), denoted by $G=(V,E,\bm{c}+\theta\bm{\tilde{c}})$. The edge costs are parametrized by the value $\theta$ and are given as the original edge costs ($\bm{c}$) plus $\theta$ times a perturbation cost vector ($\bm{\tilde{c}}$). The decision variables are denoted by $x_{jk}$ and equal 1 if edge $(v_{j},v_{k})$ is used in the solution and 0 otherwise. The parametrized SPP is then given as
$\min\{(\bm{c}+\theta\bm{\tilde{c}})^{\intercal}\bm{x}:\bm{x}\in\mathbb{X}_{SPP}\},$
where $\mathbb{X}_{SPP}$ denotes the set of incidence vectors of all paths in the graph. The full formulation can be found in Equation 13 of the appendix.
We examine how the objective value and decision variable values of the original instance are affected by parameter $\theta$. We consider the instance of SPP-$\theta$ as displayed in Figure 1 with $\theta^{0}=0$ as the present problem. By varying $\theta$, the optimal shortest $(s,t)$-path and its optimal value changes.
[…]
Figure 1: Instance of SPP-$\theta$, with in blue the optimal $s,t$-route found using Dijkstra’s algorithm for the present problem, i.e., $\theta^{0}=0$.
We sample $\theta$ uniformly on the interval $[-1,1]$ and run CLEMO on the sampled data. In Figure 2(a), we show the true dependency of the optimal value of SPP-$\theta$ and $\theta$ and the prediction of CLEMO. Figure 2(b) shows the same for the dependency of three selected decision variable values. For the predictions of all decision variables, see Figure 4 in the appendix. Both results show that CLEMO manages to generate locally accurate predictions. In Table 1, we show the accuracy and the incoherence of CLEMO and the LR benchmark. We can conclude that our method finds significantly more coherent explanations without considerably conceding accuracy.
(a) CLEMO prediction of objective value.
(b) CLEMO prediction of a subset of decision variables.
@@ 4.3 Knapsack Problem
Next, we present an extensive study on the Knapsack problem (KP). In this problem, we are given a set of items each with a corresponding value $v_{j}$ and weight $w_{j}$. The goal is to decide how much of each item should be chosen to maximize the total value while not exceeding the capacity, which w.l.o.g. we set to $1$. Formulated as a linear problem this becomes $\max\{\bm{v}^{\intercal}\bm{x}:\bm{w}^{\intercal}\bm{x}\leq 1,\bm{x}\in[0,1]^{p}\}$. We consider the parametrized KP, by setting $\bm{\theta}=(\bm{v},\bm{w})$ and use Gurobi to solve to optimality.
This experiment compares CLEMO with two benchmark explanation methods: independently fitting the components using (i) a linear regression model (LR), and (ii) a decision tree regressor (DTR) with a maximum depth of 5, and minimum samples per leaf of 50. Here, we consider 40 instances of the KP each with $p=25$ items. The 40 instances are divided over four instance types as described by Pisinger (2005): 1) uncorrelated, 2) weakly correlated, 3) strongly correlated, 4) inversely strongly correlated.
In Table 2, we see that compared to a linear regression approach the linear model found by CLEMO reduces the weighted incoherence in the objective and the constraint by more than 50% and 99% respectively, while the weighted accuracy loss increased only by roughly 20%. In Figures 5, 6, 7 and 8 in the appendix we plotted the accuracy and incoherence of each instance to strengthen our conclusion.
Table 2: Mean weighted accuracy loss and incoherence for the KP solved optimally.
@@
Method Accuracy ($\ell_{A}$) Objective
Feasible
region
Type 1 DTR 405 6.48 22.50
LR 437 0.24 5.49
CLEMO 479 0.08 0.01
Type 2 DTR 868 20.09 38.55
LR 1076 1.01 9.49
CLEMO 1203 0.38 0.03
Type 3 DTR 865 28.45 39.41
LR 1061 1.39 9.90
CLEMO 1189 0.58 0.04
Type 4 DTR 893 13.80 37.74
LR $1$ 0.70 8.94
CLEMO 1241 0.27 0.03
Besides, as datasets are randomly generated we measure the stability of explanations over resampling. For the KP, we analyze the stability of CLEMO by using 10 different randomly generated datasets, resulting in 10 surrogate models. To quantify stability we use the (normalized) standard deviation of the feature contributions of $g$ which is also used to examine the stability of LIME (Shankaranarayana & Runje, 2019). In Table 3, we consider the (normalized) standard deviation of the contribution of the top-5 most contributing, non-zero features for each component of $h(\bm{\theta})$. Besides, we examine the feature stability index (FSI), which is based on the variables stability index as presented in (Visani et al., 2022). The FSI measures how much the order in feature contribution over the resamples on average coincides. Here, it takes values between 0 and 5, where a higher FSI indicates more stable explanations. An extensive description of the FSI can be found in section A.4.2 in the appendix. From Table 3, we can conclude that the stability of CLEMO is comparable to the benchmark approaches.
Table 3: Mean stability measures over 10 instances per type of KP.
Stability
Method Std. Normalized Std. FSI
Type 1 DTR 0.18 5.36 2.02
LR 0.22 1.70 2.44
CLEMO 0.18 1.75 2.26
Type 2 DTR 0.14 2.74 3.00
LR 0.19 0.74 3.40
CLEMO 0.16 0.81 3.34
Type 3 DTR 0.14 2.77 2.91
LR 0.20 0.71 3.22
CLEMO 0.16 0.79 3.23
Type 4 DTR 0.14 2.52 2.99
LR 0.18 0.66 3.31
CLEMO 0.16 0.77 3.38

### v2

@@ 3 Methodology
. (1) The function $\bm{x}\mapsto f(\bm{x};\bm{\theta})$ is affine.   (2) The function $\bm{x}\mapsto\delta(\bm{x},\mathbb{X}(\bm{\theta}))$ is convex.
The proof of this proposition follows from applying composition rules of convex functions [7]. When the conditions in this proposition are satisfied, every local minimum of the optimization problem (10) is a global minimum. We can use first-order methods to find such a minimum if the functions are differentiable. We note that for $\delta$ as defined in (9), we have that (8) is convex in $\bm{\beta}$ if the functions $\bm{x}\mapsto\gamma_{t}(\bm{x},\bm{\theta}^{i})$ are convex in $\bm{x}$ for $t\in[T]_{1}$ since $x\mapsto\max\{0,x\}$ is convex and nondecreasing in $x$. Assuming that $\bm{\beta}$ comes from a bounded space, we can then use the subgradient algorithm with constant step size and step length and ensure convergence to an $\epsilon$-optimal point within a finite number of steps in $\mathcal{O}(1/\epsilon^{2})$ [8].
Weights. We define the weights (similarly to LIME) as the radial basis function kernel with kernel parameter $\nu$ and distance function $d$,
$w^{i}=\exp(-d(\bm{\theta}^{i},\bm{\theta}^{0})^{2}/\nu^{2}),\quad i\in[N]_{0}.$ (11)
Sampling. We recall that contrary to ML models, optimization models do not require model training per se. Therefore, $\bm{\theta}^{i}$ cannot be sampled according to the train data distribution. Depending on the context, $\bm{\theta}^{i}$ can be sampled from relevant distributions, or with pre-determined rules, e.g., discretization. Besides, we note that for some values of $\bm{\theta}^{i}$ the optimization problem might be infeasible or unbounded. We therefore ensure the generated dataset $\mathcal{D}$ contains only feasible, bounded instances of the optimization problem.
Binary Decision Variables. We opt for logistic regression to obtain interpretable surrogate models for the output components of the optimization problem that are restricted to binary values. Let $\mathcal{B}\subseteq\{f,x_{1},\ldots,x_{p}\}$ be the set of binary components. Then, for a binary component $c\in\mathcal{B}$, we consider predictors of the form $g_{c}(\bm{\theta})=\sigma(\bm{\beta^{\intercal}}_{c}\bm{\theta})$ with $\sigma:\mathbb{R}\to[0,1]$ the sigmoid function. The relative values of vector $\bm{\beta}_{c}$ then tell the user the feature importance for the probability of the component being 00 or $1$. For the binary components, we measure the prediction accuracy using the log-loss. The total accuracy loss then becomes
$\lambda_{A_{1}}\sum\nolimits_{c\in\overline{\mathcal{B}}}\|\bm{\beta^{\intercal}}_{c}\bm{\theta}^{i}-h_{c}(\bm{\theta}^{i})\|^{2}-\lambda_{A_{2}}\sum\nolimits_{c\in\mathcal{B}}h_{c}(\bm{\theta}^{i})\ln(\sigma(\bm{\beta^{\intercal}}_{c}\bm{\theta}^{i}))+(1-h_{c}(\bm{\theta}^{i}))\ln(1-\sigma(\bm{\beta^{\intercal}}_{c}\bm{\theta}^{i})),$
@@
Guaranteed Objective Coherence in the Linear Case. Assume our present problem (2) has a linear objective function, i.e., the objective is of the form $f(\bm{x};\bm{\theta})=\bm{\hat{c}}^{\intercal}\bm{x}$, and assume that only the feasible region is sensitive, i.e., $\bm{\hat{c}}$ remains fixed. In this case, we can fit $p+1$ independent linear models for each component in $c\in\{f,x_{1},\ldots,x_{p}\}$ by solving the classical weighted mean-square problem
$\min_{\bm{\beta}_{c}}\ \sum\nolimits_{i=0}^{N}w^{i}\|\bm{\beta}_{c}^{\intercal}\bm{\theta}^{i}-h_{c}(\bm{\theta}^{i})\|^{2}.$
If the minimizers are unique, the corresponding linear predictors provably fulfill the coherence condition (3), i.e., in this case we do not need to apply the regularizer $R_{C}$ to achieve coherence condition (3). However, it may happen that condition (4) is violated as the example from the introduction shows. A proof of the latter coherence statement can be found in Section A.3 (Theorem A.1).
@@ 4 Experiments
In this section, we present three experiments. Each experiment considers a distinct optimization model and solver. The first experiment is used as a proof of concept, where we will show that CLEMO approximates the optimal decision and objective value function well for an instance of the Shortest Path Problem (SPP) with a single sensitive parameter. For a second experiment, in an extensive study, we consider exact solutions of various instances of the Knapsack Problem (KP). We compare the quality of explanations found by CLEMO to benchmarks by analyzing local accuracy, coherence, and stability of the found explanations when subjected to resampling. Lastly, we generate explanations for the Google OR-Tools heuristic [13] applied to an instance of the Capacitated Vehicle Routing Problem (CVRP). The code of our experiments can be found at https://github.com/daanotto/CLEMO. All experiments are done on a computer with a 13th Gen Intel(R) Core(TM) i7-1355U 1.70 GHz processor and 64 GB of installed RAM.
Setup. Unless stated otherwise, all upcoming experiments use the following setup. Given an optimization problem for a given parameter vector $\bm{\theta}^{0}$, we create a training data set $\mathcal{D}$ of size 1000 by sampling $\bm{\theta}^{i}\sim\mathcal{N}(\bm{\theta}^{0},0.2\bm{\theta}^{0})$. The sample’s proximity weights $w^{i}$ are determined using the rbf kernel (11) with Euclidean distance and parameter $\nu$ equal to the mean distance to $\bm{\theta}^{0}$ over the training data set $\mathcal{D}$.
As a benchmark for CLEMO, we consider generating explanations with the LIME-type method described in Section 3. We solve problem (5) where we fit logistic regression models for all binary output components and linear models for all other output components without any complexity regularization. We refer to this benchmark as LR.
For CLEMO we use the loss function stated in (10), where $\ell_{A}$ is given as in (3) and $R_{C}$ is given as in (8) with $\delta$ defined as in (9). This way we can compare CLEMO to the benchmark on local accuracy (3) and incoherence (8). In CLEMO, each term of the total loss function is weighted with hyperparameters $\lambda_{A_{1}},\lambda_{A_{2}},\lambda_{C_{1}},$ and $\lambda_{C_{2}}$ as
$\lambda_{A_{1}}\ell_{A_{1}}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))+\lambda_{A_{2}}\ell_{A_{2}}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))+\lambda_{C_{1}}R_{C_{1}}(g(\bm{\theta}^{i}))+\lambda_{C_{2}}R_{C_{2}}(g(\bm{\theta}^{i})).$
@@ 4.1 Shortest Path Problem
As a first experiment, we explain an instance of the Shortest Path where possible cost-changes depend on one single parameter. An instance of the SPP is given by a connected graph $G=(V,E,\bm{c})$, with nodes $V$, edges $E$ and edge-costs $\bm{c}$, and specified start and terminal nodes $s,t\in V$. The objective is to find a path between $s$ and $t$ of minimum costs. Several methods exist for solving the SPP in polynomial time [14]. Here, we use Dijkstra’s algorithm.
We study the parametric version of the shortest path problem (SPP-$\theta$), denoted by $G=(V,E,\bm{c}+\theta\bm{\tilde{c}})$. The edge costs are parametrized by the value $\theta$ and are given as the original edge costs ($\bm{c}$) plus $\theta$ times a perturbation cost vector ($\bm{\tilde{c}}$). The decision variables are denoted by $x_{jk}$ and equal 1, if edge $(v_{j},v_{k})$ is used in the solution and 0, otherwise. The parametrized SPP is then given as $\min\{(\bm{c}+\theta\bm{\tilde{c}})^{\intercal}\bm{x}:\bm{x}\in\mathbb{X}_{SPP}\}$, where $\mathbb{X}_{SPP}$ denotes the set of incidence vectors of all paths in the graph. The full formulation can be found in Equation 12 of the appendix. We examine how the objective value and decision variable values of the original instance are affected by parameter $\theta$. We consider the instance of SPP-$\theta$ as displayed in Figure 1(a) with $\theta^{0}=0$ as the present problem. By varying $\theta$, the optimal shortest $(s,t)$-path and its optimal value changes.
(a) Instance of SPP-$\theta$, with in blue the optimal $s,t$-route.
(b) CLEMO prediction of objective value.
(c) CLEMO prediction of decision variable $x_{(s,a)}$.
We sample $\theta$ uniformly on the interval $[-1,1]$ and run CLEMO on the sampled data. In Figure 1(b), we show the true dependency of the optimal value of SPP-$\theta$ and $\theta$ and the prediction of CLEMO. Figure 1(c) shows the same for the dependency of three selected decision variable values. For the predictions of all decision variables, see Figure 4 in the appendix. Both results show that CLEMO manages to generate locally accurate predictions. In Table 1, we show the accuracy and the incoherence of CLEMO and the LR benchmark. We can conclude that our method finds significantly more coherent explanations without considerably conceding accuracy.
@@
Feasible region
@@ 4.2 Knapsack Problem
Next, we present an extensive study on the Knapsack problem (KP). In this problem, we are given a set of items each with a corresponding value $v_{j}$ and weight $w_{j}$. The goal is to decide how much of each item should be chosen to maximize the total value while not exceeding the capacity, which w.l.o.g. we set to $1$. Formulated as a linear problem this becomes $\max\{\bm{v}^{\intercal}\bm{x}:\bm{w}^{\intercal}\bm{x}\leq 1,\bm{x}\in[0,1]^{p}\}$. We consider the parametrized KP, by setting $\bm{\theta}=(\bm{v},\bm{w})$ and use Gurobi [17] to solve to optimality. This experiment compares CLEMO with two benchmark explanation methods: independently fitting the components using (i) a linear regression model (LR), and (ii) a decision tree regressor (DTR) with a maximum depth of 5, and minimum samples per leaf of 50. Here, we consider 40 instances of the KP each with $p=25$ items. The 40 instances are divided over four instance types as described by [28]: 1) uncorrelated, 2) weakly correlated, 3) strongly correlated, 4) inversely strongly correlated.
In Table 2, we see that compared to a linear regression approach the linear model found by CLEMO reduces the weighted incoherence in the objective and the constraint by more than 50% and 99% respectively, while the weighted accuracy loss increased only by roughly 20%. In Figures 5, 6, 7 and 8 in the appendix, we plotted the accuracy and incoherence of each instance to strengthen our conclusion.
Table 2: Mean $(\mu)$ and standard deviation ($\sigma$) of weighted accuracy loss and incoherence for the KP solved optimally. On the right, the mean stability measures over 10 instances per type of KP.
@@
Accuracy ($\ell_{A}$) Objective Feasible region Stability
Method $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ Std. Normalized Std. FSI
Type 1 DTR 405 97.4 6.48 2.55 22.50 4.99 0.18 5.36 2.02
LR 437 140 0.24 0.15 5.49 1.51 0.22 1.70 2.44
CLEMO 479 157 0.08 0.06 0.01 0.004 0.18 1.75 2.26
Type 2 DTR 868 67.4 20.09 2.27 38.55 2.03 0.14 2.74 3.00
LR 1076 99.4 1.01 0.12 9.49 0.68 0.19 0.74 3.40
CLEMO 1203 113 0.38 0.06 0.03 0.01 0.16 0.81 3.34
Type 3 DTR 865 66.8 28.45 3.36 39.41 2.02 0.14 2.77 2.91
LR 1061 97.8 1.39 0.16 9.90 0.64 0.20 0.71 3.22
CLEMO 1189 113 0.58 0.07 0.04 0.01 0.16 0.79 3.23
Type 4 DTR 893 56.5 13.80 2.16 37.74 2.35 0.14 2.52 2.99
LR $1$ 77.6 0.70 0.11 8.94 0.74 0.18 0.66 3.31
CLEMO 1241 87.56 0.27 0.06 0.03 0.01 0.16 0.77 3.38
Table 3: Runtime of different explanation approaches for various sizes of KP.
KP Runtime (s)
#items #features DTR LR CLEMO
5 10 0.0189 0.0031 8.54
10 20 0.0396 0.0790 50.0
20 40 0.183 0.0830 413
40 80 0.683 0.316 $>1000$
Figure 2: Convergence of CLEMO over SLSQP iterations for different sizes of KP.
Besides, as datasets are randomly generated, we measure the stability of explanations over resampling. For the KP, we analyze the stability of CLEMO by using 10 different randomly generated datasets, resulting in 10 surrogate models. To quantify stability we use the (normalized) standard deviation of the feature contributions of $g$ which is also used to examine the stability of LIME [33]. In Table 2, we consider the (normalized) standard deviation of the contribution of the top-5 most contributing, nonzero features for each component of $h(\bm{\theta})$. Besides, we examine the feature stability index (FSI), which is based on the variables stability index as presented in [35]. The FSI measures how much the order in feature contribution over the resamples on average coincides. Here, it takes values between 0 and 5, where a higher FSI indicates more stable explanations. An extensive description of the FSI can be found in section A.4.2 in the appendix. From Table 2, we can conclude that the stability of CLEMO is comparable to the benchmark approaches.
Runtime. To compare the runtime of CLEMO to benchmark methods, we analyze the runtime on Knapsack problems with 5, 10, 20, and 40 items. In Table 3, we see CLEMO takes longer to find an explanation. Looking at Figure 2, we observe that CLEMO efficiently converges to a solution. Hence, early stopping could reduce runtimes while still ensuring more coherent explanations.